A business has a current cash flow of sh 400,000. Currently, its cost of capital is estimated at 10%. Required a) Assuming it will generate these cash flows every year till infinity, how much is it worth today. b) Assume the cash flows will grow at 8% in the first 4 years and stabilize at 5% from year five on wards. Calculate the value of the business today if return required by investors today is 12%
Answer #1

Answer:

A)

Given Cash flow C=400000

Rate r=10%

So value of Company =C/r=400000/10%=4000000

B)

Given

Current Cash flow in C0=400000

Growth rate in first four years G=8%

Cash flow in first year C1=400000*(1+8%)=432000

Cash flow in Second year C2=400000*(1+8%)^2=466560

Cash flow in Three year C3=400000*(1+8%)^3=503884

Cash flow in Four year C4=400000*(1+8%)^4=544195.6

After that growth rate in cash flow is g =5%

Rate r=12%

So Value of Company =C0+C1/(1+r) +C2/(1+r)^2 + C3/(1+r)^3 + C4/(1+r)^4 + C4*(1+g)/(r-g)

So Value of Company =400000+432000/(1+12%) +466560/(1+12%)^2 + 503884/(1+12%)^3 + 544195.6/(1+12%)^4 + 544195.6*(1+5%)/(12%-5%)=10025088.2